# Invariant Subspace Decomposition

### v1

@@ D.1 Proof of Theorem 1
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ and define $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ as in (10). By Assumption 2, $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ form an orthogonal partition w.r.t. $(X_{t})_{t\in\mathbb{N}}$ and $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant on $\mathbb{N}$. Therefore, using Lemma 4, we get for all $t\in\mathbb{N}$ that
@@
Moreover, Assumption 1 ensures that the invariant and residual subspaces can be uniquely identified from an arbitrary irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$, and therefore all the above results do not depend on the considered partition. This concludes the proof of Theorem 1. ∎
D.2 Proof of Proposition 1
@@
For all $j\in\{1,\dots,q^{U}_{\max}\}$, let $U^{S_{j}}$ denote the submatrix of $U$ formed by the columns indexed by $S_{j}$. Then, the orthogonal projection matrix onto the subspace $\mathcal{S}_{j}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$. It follows for all $t\in[n]$ that
@@
where the last equality holds since $(U^{S_{i}})^{\top}\Sigma_{t}U^{S_{j}}$ is the $(i,j)$-th (off-diagonal) block of $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$, which is zero due to the block diagonal structure of $\tilde{\Sigma}_{t}$. Irreducibility of the orthogonal partition follows from the irreducibility of the joint block diagonal decomposition. ∎
D.3 Proof of Proposition 2
@@
(i) First, observe that, by Lemma 6, $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ and that, by Lemma 5, $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant on $[n]$. Then, by definition of $\beta^{\operatorname{inv}}$ and by Lemma 8 we get for all $t\in[n]$ that
@@
where the third equality uses $\beta^{\operatorname{inv}}=\Pi_{\mathcal{S}^{\operatorname{inv}}}\gamma_{0,t}$ by Proposition 2 $(i)$ and the last equality follows from the fact that $\mathcal{S}^{\operatorname{inv}}$, $\mathcal{S}^{\operatorname{res}}$ are an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ by Lemma 6.
(iii) Let $\mathcal{B}^{n}$ denote the set of all time-invariant parameters over $[n]$. By assumption, we have that $\mathcal{B}^{n}\subseteq\mathcal{S}^{\operatorname{inv}}$. Moreover, by point $(ii)$, we have that $\beta^{\operatorname{inv}}\in\mathcal{B}^{n}$. Therefore, by definition of $\beta^{\operatorname{inv}}$ we obtain that $\beta^{\operatorname{inv}}=\operatorname*{arg\,max}_{\beta\in\mathcal{B}^{n}}\overline{\operatorname{\Delta Var}}(\beta)$. This completes the proof of Proposition 2. ∎
D.4 Proof of Theorem 2
@@ D.5 Proof of Theorem 3
where $\mathbf{\Sigma}\coloneqq\mathbb{E}[\frac{1}{n}\mathbf{X}^{\top}\mathbf{X}]$. Using now that, by assumption, $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal partition w.r.t. $(X_{t})_{t\in\mathbb{N}}$, let $U\in\mathbb{R}^{p\times p}$ be defined for such a partition as in Lemma 2 and let $U^{\operatorname{inv}}\in\mathbb{R}^{p\times\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})}$ be the submatrix of $U$ whose columns form an orthonormal basis for $\mathcal{S}^{\operatorname{inv}}$, such that $\Pi_{\mathcal{S}^{\operatorname{inv}}}=U^{\operatorname{inv}}(U^{\operatorname{inv}})^{\top}$. Moreover, let $\tilde{\mathbf{\Sigma}}^{\operatorname{inv}}\coloneqq(U^{\operatorname{inv}})^{\top}\mathbf{\Sigma}U^{\operatorname{inv}}$ and $\tilde{\Sigma}_{t^{*}}^{\operatorname{inv}}\coloneqq(U^{\operatorname{inv}})^{\top}\Sigma_{t^{*}}U^{\operatorname{inv}}$ denote, as in Lemma 2, the diagonal block corresponding to $\mathcal{S}^{\operatorname{inv}}$ of the block diagonal matrices $U^{\top}\mathbf{\Sigma}U$ and $U^{\top}\Sigma_{t^{*}}U$, respectively. By the properties of orthogonal matrices, $(\tilde{\mathbf{\Sigma}}^{\operatorname{inv}})^{-1}=(U^{\operatorname{inv}})^{\top}\mathbf{\Sigma}^{-1}U^{\operatorname{inv}}$. For all $i\in\{1,\dots,\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})\}$, let $\lambda_{i}$ denote the $i$-th eigenvalue in decreasing order. Then, we can further express the above lower bound as
@@ D.6 Proof of Proposition 3
(ii) By definition of $\beta^{\operatorname{inv}}$ and using point (i) of this proposition and that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ form a partition w.r.t. $(X_{t})_{t\in[n]}$, we obtain that, for all $t\in[n]$,
@@
(iii) The claim follows from the definition of $\delta^{\operatorname{res}}_{t}$, from the fact that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ forms a partition w.r.t. $(X_{t})_{t\in[n]}$ and from point (i) of this proposition. ∎

### v2

@@ E.1 Proof of Theorem 1
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition and define $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ as in (11). By Assumption 2, $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ form an orthogonal and $(X_{t})_{t\in\mathbb{N}}$-decorrelating partition and $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant on $\mathbb{N}$. Therefore, using Lemma 1 and Lemma 2, we get for all $t\in\mathbb{N}$ that
@@
Moreover, Assumption 1 ensures that the invariant and residual subspaces can be uniquely identified from an arbitrary irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition, and therefore all the above results do not depend on the considered partition. This concludes the proof of Theorem 1. ∎
E.2 Proof of Proposition 1
@@
(i) For all $j\in\{1,\dots,q^{U}_{\max}\}$, let $U^{S_{j}}$ denote the submatrix of $U$ formed by the columns indexed by $S_{j}$. Then, the orthogonal projection matrix onto the subspace $\mathcal{S}_{j}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$. It follows for all $t\in[n]$ that
@@
where the last equality holds since $(U^{S_{i}})^{\top}\Sigma_{t}U^{S_{j}}$ is the $(i,j)$-th (off-diagonal) block of $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$, which is zero due to the block diagonal structure of $\tilde{\Sigma}_{t}$. Irreducibility of the orthogonal partition follows from the irreducibility of the joint block diagonal decomposition.
(ii) The statement follows from Lemma 4, taking $\mathcal{N}=[n]$. Irreducibility of the joint block diagonalizer follows from the irreducibility of the orthogonal partition. ∎
E.3 Proof of Proposition 2
@@
(i) First, observe that, by Lemma 6, $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition and that, by Lemma 7, $\mathcal{S}^{\operatorname{inv}}$ is opt-invariant on $[n]$. Then, by definition of $\beta^{\operatorname{inv}}$ and by Lemma 9 we get for all $t\in[n]$ that
@@
where the third equality uses $\beta^{\operatorname{inv}}=\Pi_{\mathcal{S}^{\operatorname{inv}}}\gamma_{0,t}$ by Proposition 2 $(i)$ and the last equality follows from the fact that $\mathcal{S}^{\operatorname{inv}}$, $\mathcal{S}^{\operatorname{res}}$ are an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition by Lemma 6.
(iii) Let $\mathcal{B}^{n}$ denote the set of all time-invariant parameters over $[n]$. By assumption, we have that $\mathcal{B}^{n}\subseteq\mathcal{S}^{\operatorname{inv}}$. Moreover, by point $(ii)$, we have that $\beta^{\operatorname{inv}}\in\mathcal{B}^{n}$. Therefore, by definition of $\beta^{\operatorname{inv}}$ we obtain that $\beta^{\operatorname{inv}}=\operatorname*{arg\,max}_{\beta\in\mathcal{B}^{n}}\overline{\operatorname{\Delta Var}}(\beta)$.
This completes the proof of Proposition 2. ∎
E.4 Proof of Theorem 2
@@ E.5 Proof of Theorem 3
where $\mathbf{\Sigma}\coloneqq\mathbb{E}[\frac{1}{n}\mathbf{X}^{\top}\mathbf{X}]$. Using now that, by assumption, $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal and $(X_{t})_{t\in\mathbb{N}}$-decorrelating partition, let $U\in\mathbb{R}^{p\times p}$ be defined for such a partition as in Lemma 4 and let $U^{\operatorname{inv}}\in\mathbb{R}^{p\times\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})}$ be the submatrix of $U$ whose columns form an orthonormal basis for $\mathcal{S}^{\operatorname{inv}}$, such that $\Pi_{\mathcal{S}^{\operatorname{inv}}}=U^{\operatorname{inv}}(U^{\operatorname{inv}})^{\top}$. Moreover, let $\tilde{\mathbf{\Sigma}}^{\operatorname{inv}}\coloneqq(U^{\operatorname{inv}})^{\top}\mathbf{\Sigma}U^{\operatorname{inv}}$ and $\tilde{\Sigma}_{t^{*}}^{\operatorname{inv}}\coloneqq(U^{\operatorname{inv}})^{\top}\Sigma_{t^{*}}U^{\operatorname{inv}}$ denote, as in Lemma 4, the diagonal block corresponding to $\mathcal{S}^{\operatorname{inv}}$ of the block diagonal matrices $U^{\top}\mathbf{\Sigma}U$ and $U^{\top}\Sigma_{t^{*}}U$, respectively. By the properties of orthogonal matrices, $(\tilde{\mathbf{\Sigma}}^{\operatorname{inv}})^{-1}=(U^{\operatorname{inv}})^{\top}\mathbf{\Sigma}^{-1}U^{\operatorname{inv}}$. For all $i\in\{1,\dots,\operatorname{dim}(\mathcal{S}^{\operatorname{inv}})\}$, let $\lambda_{i}$ denote the $i$-th eigenvalue in decreasing order. Then, we can further express the above lower bound as
@@ E.6 Proof of Proposition 3
(ii) By definition of $\beta^{\operatorname{inv}}$ and using point (i) of this proposition and that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ form a $(X_{t})_{t\in[n]}$-decorrelating partition, we obtain that, for all $t\in[n]$,
@@
(iii) The claim follows from the definition of $\delta^{\operatorname{res}}_{t}$, from the fact that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ forms a $(X_{t})_{t\in[n]}$-decorrelating partition and from point (i) of this proposition. ∎